OpenStudy (anonymous):

One number is 5 more than another number. If you add 8 to 2 times the first number, the result is 3 times the second number.

Directrix (directrix):

Let x be one number Then, x + 5 is the "another" 2 x + 8 = 3 (x +5) 2x +8 = 3x + 15 -x = 7 x = -7 Other number is - 7 + 5 = -2